OpenStudy (anonymous):

simplify -2[-13-(-52-11)]=

OpenStudy (anonymous):

-2 * (-13 + 63) = -2* 50 = -100

OpenStudy (anonymous):

so -100 is the answer?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

first work out inner brackets = -63 then as there is a '-' outside the brackets this makes +63

OpenStudy (anonymous):

I'm guessing this is a question about the order of operations. Start with the stuff in the deepest bracket first and work your way out.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

-100